A firm has total revenue of $50 million ,fixed cost of $60 million,and variable cost of $40 million.What does the firm do (a) in the short run?(b) in the long run?
1
Expert's answer
2019-03-21T12:42:52-0400


The company needs to increase production levels to achieve a break-even level.

The break-even is sales volume at which the profit of the entrepreneur is zero.

The break-even point in monetary terms, is also called the “profitability threshold”, is calculated using the following formula:

BEP=FC/KMR.At the existing production level, the break-even level can be calculated as follows:


KMR=MR/TR

MR=TR-VC=50-40=10

KMR=10/50=0.2

BEP=60/0.2=300 mln



In long run, this firm should rises its price, or produce volume over the BEP, or exit the market.